Exabytes - Schedule Hard Drive Upgrade on Leo (202.75.42.86)

Date : 9 September 2009
Time : 6.00 PM GMT+ 0800

Maintenance Progress:
------------------------
The rsync process has been completed and currently all services running fine in LEO server. Kindly email us immediately if there is any issue occurred after this maintenance for our further assistance.

======================================================================================================================

Date : 9 September 2009
Time : 11:00 AM GMT+ 0800

Maintenance Progress:
------------------------
The rsync process is 60% completed.

======================================================================================================================

Date : 9 September 2009
Time : 9:50 AM GMT+ 0800

Maintenance Progress:
------------------------
The rsync process is 30% completed.

======================================================================================================================

Date : 9 September 2009
Time : 8.30 AM GMT+ 0800

Maintenance Progress:
------------------------
LEO HDD upgrade is completed. All services resume.
HDDs are running mirror rsync background.

======================================================================================================================


Date : 9 September 2009
Time : 6.00 AM GMT+ 0800
Expected ETA : 2 Hours

Maintenance Progress:
------------------------
LEO image restoration still in progress. At this moment the progress is still 50%.
======================================================================================================================

Date : 9 September 2009
Time : 4.30 AM GMT+ 0800
Expected ETA : 4 Hours

Maintenance Progress:
------------------------
The backup image of LEO server has been generated and the old HDD's has been replaced with a new HDD's. At this moment, we are in the midst of restoring back the image to the server.

======================================================================================================================

Date : 9 September 2009
Time : 2.00 AM GMT+ 0800
Expected ETA : 6 Hours

Maintenance Progress:
------------------------
At this moment, the actual image of LEO server currently still being compress and transferred to external hard drive. Current progress already reach 60% of creating a backup image.
We will update this announcement for next progress.

======================================================================================================================

Date : 8 September 2009
Time : 11.25PM GMT+ 0800
Expected ETA : 6 Hours

Maintenance Progress:
------------------------
At this moment, the actual image of LEO server currently being compress and transferred to external hard drive. We estimate this backup phase shall be completed within 2hrs time. Once the backup image is created, we will continue with 2nd phase which is hardware upgrade (replace existing hdd's with new hdd's). We will update this announcement for next progress.

======================================================================================================================

Date : 8 September 2009
Time : 10.00PM GMT+ 0800
Expected ETA : 5 Hours

Maintenance Details:
------------------------
We will be upgrading the existing hard disk for the Windows Web Server - Leo from 200GB to 500GB as to support the continuous growth of disk usage for existing clients on the server.

Maintenance Effect:
------------------------
As we need to synchronize the existing data to the new hard disk and making sure all the data is up-to-date, we will be stopping the all the services such as FTP, Web and MySQL on the server side. We estimate the entire process will completes in 4 hours time. Mail service operate as usual as it is hosted